OpenStudy (anonymous):

a person invested $60000 into two accounts that pays 8% and 6% simple interest annually, respectively. Find the amount invested into the two accounts if the total interest earned after one year is $ 3600

OpenStudy (anonymous):

\[.08\times x + .06\times (60000-x)=3600\] solve for x

OpenStudy (anonymous):

that is, let x = amount invested at 8% and so 6000-x is amount invested at 6%. do you know how to solve this for x?

OpenStudy (anonymous):

turns out it was all invested at 6% because 6% of $60,000 = $3600 exactly

OpenStudy (anonymous):

let sum invested at 8% = A interest1 = PRT/100 = (A x 8 x 1)/100 = 8A/100 sum invested at 6% = 60000 - A interest2 = PRT/100 = (60000 - A) x 6 x 1/100 = (360000 - 6A)/100 total interest = interest1 + interest 2 = 8A/100 + (36000 - 6A)/100 = (8A - 6A + 36000)/100 = (2A + 36000)/100 ATQ total interest is 3600 so 2A + 360000 ------------ = 3600 100 2A + 360000 = 3600 x 100 2A = 360000 - 360000 2A = 0 so A = 0/2 = 0 This means that sum invested at 8% = 0 and sum invested at 6% = 60000